OpenStudy (anonymous):

Is this the same as the last question? @Hero Find the zeros of the polynomial function. f(x) = x2 – 4x – 60 A. –6, 10 B. –10, 6 C. –14, 5 D. –5, 14

OpenStudy (shamil98):

(x+6)(x-10) so -6 , and 10.

OpenStudy (shamil98):

When factoring out the equation into that form. (x+6)(x-10) = 0 x - 10 = 0 x = 10 or x + 6 = 0 x= -6